How do you write the slope-intercept form of the equation for the line passing yhrough the given pair of points. I used this formula m=y2-y1/x2-x1 with y2=7/5,y1=4/5,x2=1/5,x1=-1/5
Answer by Alan3354(69443)

How do you write the slope-intercept form of the equation for the line passing yhrough the given pair of points. I used this formula m=y2-y1/x2-x1 with y2=7/5,y1=4/5,x2=1/5,x1=-1/5
--------------
You didn't give the points, but the appear to be
(-1/5,4/5) and (1/5,7/5)
Slope m = diffy/diffx = (3/5 // 2/5) = 3/2
Then use y = mx + b with either point to find b
7/5 = (3/2)*(1/5) + b
b = 11/10
------------
y = (3/2)x + 11/10